# Aeon: High-Performance Neuro-Symbolic Memory Management for Long-Horizon LLM Agents

### v1

@@ 5.1 Hardware Environment
We evaluate Aeon on a single Apple M4 Max workstation, representing a high-end consumer-grade ARM64 platform. The system specification is as follows:
@@ 5.2 Datasets
We evaluate Aeon using synthetic datasets designed to stress-test the Atlas index under controlled conditions.
@@ 5.2.1 Synthetic Atlas: Dense Forest
We generate a “Dense Forest” of synthetic vectors to simulate a large-scale knowledge base. Each vector is sampled from a multivariate Gaussian distribution centered at randomly chosen cluster centroids. Dataset sizes are varied across three orders of magnitude:
@@ 5.2.2 Workload Traces
We define two distinct workload profiles to simulate real-world query patterns:
@@ 5.3 Baselines
We compare Aeon against two baseline systems representing the spectrum from naive to state-of-the-art approaches:
@@
We evaluate two configurations of Aeon to isolate the contribution of the Semantic Lookaside Buffer (SLB):
@@ 5.4 Metrics
We report the following metrics to provide a comprehensive performance profile:
@@
QPS (Queries Per Second): Throughput measured under sustained load. We report peak QPS achieved when queries are issued in a tight loop with no artificial delays.
@@
All latency measurements are taken using std::chrono::high_resolution_clock with nanosecond precision. We exclude the first 100 queries from each run to allow the system to reach a steady state (warm caches, JIT compilation for Python components).
@@ 6.1 Micro-Benchmarks: The Kernel Speed
Result. The AVX-512 SIMD kernel achieves approximately 50 nanoseconds per 768-dimensional vector comparison. This represents a dramatic improvement over alternative implementations: 20 faster than an equivalent scalar C++ loop, and nearly 2000 faster than a Python loop using NumPy element-wise operations.
Analysis. These gains stem from three synergistic optimizations. First, the AVX-512 instruction set processes 16 single-precision floats per cycle, amortizing instruction fetch and decode overhead. Second, our 4 loop unrolling strategy maximizes instruction-level parallelism by keeping multiple vector registers in flight simultaneously. Third, explicit prefetch hints (_mm_prefetch) ensure that subsequent cache lines are loaded before they are needed, eliminating memory stalls.
The practical implication is substantial: at 50ns per comparison, the math kernel alone can evaluate 20 million vector pairs per second on a single core. This raw throughput is the engine that powers the SLB’s brute-force strategy—scanning 64 cached nodes takes only 3.2 microseconds, well within L1 cache access latency budgets.
@@ 6.2 Macro-Benchmarks: The SLB Impact
Hit latency: 0.05ms (50s). The SLB provides an entry point within the target neighborhood, requiring only 1–2 additional hops to reach the optimal node.
Miss latency: 2.50ms. The search falls back to a full root-to-leaf traversal of the Atlas tree, incurring the full cost of main memory access.
@@
Comparison. The HNSW baseline exhibits constant latency of approximately 1.5ms regardless of query locality, as it cannot exploit temporal correlation between queries. Consequently, Aeon outperforms HNSW by a factor of 3 on average, and by 30 on cache hits. This advantage grows with workload locality: in highly focused conversational sessions, hit rates approach 95%, reducing effective latency to under 0.2ms.
@@ 6.3 Scalability: From 10K to 1M Nodes
Result. Flat (brute-force) search exhibits linear scaling: latency grows proportionally with database size, reaching over 100ms at nodes. In stark contrast, Aeon Atlas demonstrates logarithmic scaling: latency increases from 0.8ms to only 2.5ms across the same range—a factor of 40 improvement at enterprise scale.
Analysis. This behavior is a direct consequence of the B+ tree structure underlying the Atlas. Each level of the tree partitions the search space by a branching factor of , yielding complexity for tree traversal. At nodes, the tree depth is only levels, requiring at most 4 node comparisons to reach any leaf.
@@ 6.4 The Zero-Copy Overhead
Result. Transferring 10MB of vector data from C++ to Python incurs only 2 microseconds of latency when using zero-copy shared memory (pointer passing via mmap). In contrast, traditional serialization approaches impose catastrophic overhead:
JSON serialization: 50ms (25,000 slower)
Pickle serialization: 35ms (17,500 slower)
Analysis. The zero-copy design ensures that Python can access search results without any data movement. The 2s overhead consists solely of constructing the NumPy array header and validating memory alignment—no actual bytes are copied. This is achieved through the nanobind library’s ability to expose raw memory pointers as read-only NumPy arrays.
The practical implication is profound: the Python Shell can orchestrate complex reasoning chains—invoking the Atlas dozens of times per conversation turn—without incurring serialization penalties. A system using JSON-based IPC would spend more time marshalling data than performing actual computation. Aeon’s architecture eliminates this bottleneck entirely, enabling true real-time performance for agentic applications.
@@ 6.5 Summary
Cache Efficiency: The SLB achieves 85%+ hit rates under realistic workloads, reducing average latency to 0.42ms—3 faster than HNSW.
@@
Scalability: Logarithmic scaling ensures sub-3ms latency even at 1 million nodes.
@@
Integration Overhead: Zero-copy memory sharing eliminates serialization costs, sustaining 2s cross-language transfer.
@@ 7 Conclusion
This paper introduced Aeon, a Cognitive Operating System designed to address the fundamental limitations of stateless Large Language Models in long-horizon agentic contexts. We argued that the prevailing view of LLM memory as a simple retrieval problem is insufficient; instead, it must be treated as an active resource management task, governed by principles analogous to those found in classical operating system kernels. By formalizing semantic memory management, we demonstrated that the chaotic, probabilistic nature of vector search can be transformed into a deterministic, navigable process.
@@ 7.1 Key Contributions
Our work makes three primary contributions. First, we presented an architecture built upon the Core-Shell Zero-Copy model. The tight integration of a high-performance C++23 kernel (the Atlas) with a flexible Python reasoning layer (the Trace) via nanobind and shared memory proves viable for achieving both the latency requirements of real-time interaction and the expressiveness needed for complex neuro-symbolic reasoning. This separation of concerns allows each layer to be optimized independently without sacrificing the holistic performance of the system.
Second, we demonstrated that the Semantic Lookaside Buffer (SLB) can reduce effective retrieval latency by an order of magnitude for conversational workloads. By exploiting the inherent “semantic inertia” of human dialogue—the empirical observation that consecutive topics are highly correlated—we pre-position semantically likely nodes in a small, L1/L2 cache-resident structure. Our evaluation showed consistent hit rates exceeding 85% under realistic access patterns, validating the core hypothesis that semantic locality is a predictable and exploitable property.
Third, the Trace graph provides a level of interpretability and control that is absent in purely neural approaches. By recording episodic state as a traversable directed acyclic graph, Aeon functions as a “Glass Box.” Every decision can be traced back through its causal lineage. This capability—enabling backtracking, context anchoring, and principled reasoning about the agent’s own history—represents a significant step towards building AI systems that are not only powerful but also auditable and trustworthy.
@@ 7.2 Limitations
We acknowledge several limitations in the current design. The Atlas currently relies on static embedding models (e.g., BERT-based encoders). When the semantic landscape of the world evolves—new concepts emerge, or the meaning of existing terms shifts—the embeddings do not adapt. While the Delta Buffer provides a mechanism for ingesting new knowledge without a full rebuild, the underlying semantic geometry remains fixed. Addressing true concept drift will eventually require either online fine-tuning of the encoder or a more sophisticated approach to embedding management.
Furthermore, Aeon is presently a single-modality system, operating exclusively on text. Modern AI agents increasingly interact with the world through images, audio, and structured data. A truly general Cognitive OS must eventually support multimodal vector representations within the same spatial index, enabling unified retrieval across diverse input streams—an avenue we leave for future investigation.
@@ 7.3 Future Work
Aeon represents a first step toward a new paradigm in AI memory management, and several exciting directions remain. We envision future work exploring Multi-Tenancy and Hardware-Enforced Isolation. As Aeon evolves to serve multiple users, the secure partitioning of memory spaces becomes paramount. We are investigating the use of hardware enclaves, such as Intel SGX or ARM CCA, to provide cryptographic guarantees of data isolation at the memory level.
We also propose a “Dreaming” Process—an offline background task that activates during idle periods. This process would perform garbage collection on the Atlas, defragmenting the spatial index, and consolidate the verbose Trace into compressed, long-term episodic summaries. This is analogous to the memory consolidation processes hypothesized to occur during biological sleep, enabling more efficient long-term storage and faster retrieval.
Finally, we aim to integrate Neuro-Symbolic Reasoning more deeply into the architecture. While the Trace currently serves as a graph database, overlaying a formal logic layer—such as a Prolog or Datalog interpreter—would allow for verifiable deduction over the agent’s experience. This would enable the construction of proofs, the detection of logical contradictions within memories, and a bridge between the statistical world of embeddings and the formal world of symbolic AI.
@@ 7.4 Closing Remarks
We believe the path forward for AI agents lies not in ever-larger context windows, but in smarter, more structured memory. Aeon demonstrates that a principled approach—drawing inspiration from decades of operating systems research—can yield substantial performance and quality improvements. We hope this work contributes a useful foundation for researchers and practitioners seeking to build the next generation of persistent, coherent, and interpretable AI agents.

### v2

@@ 5.1 Hardware Environment
Aeon is evaluated on a single Apple M4 Max workstation, representing a high-end consumer-grade ARM64 platform. The system specification is as follows:
@@ 5.2 Datasets
Aeon is evaluated using synthetic datasets designed to stress-test the Atlas index under controlled conditions.
@@ 5.2.1 Synthetic Atlas: Dense Forest
A “Dense Forest” of synthetic vectors is generated to simulate a large-scale knowledge base. Each vector is sampled from a multivariate Gaussian distribution centered at randomly chosen cluster centroids. Dataset sizes are varied across three orders of magnitude:
@@ 5.2.2 Workload Traces
Two distinct workload profiles are defined to simulate real-world query patterns:
@@ 5.3 Baselines
Aeon is compared against two baseline systems representing the spectrum from naive to state-of-the-art approaches:
@@
Two configurations of Aeon are evaluated to isolate the contribution of the Semantic Lookaside Buffer (SLB):
@@ 5.4 Metrics
The following metrics are reported to provide a comprehensive performance profile:
@@
QPS (Queries Per Second): Throughput measured under sustained load. Peak QPS is reported when queries are issued in a tight loop with no artificial delays.
@@
All latency measurements are taken using std::chrono::high_resolution_clock with nanosecond precision. The first 100 queries from each run are excluded to allow the system to reach a steady state (warm caches, JIT compilation for Python components).
@@ 6.1 Micro-Benchmarks: The Kernel Speed
Result. The SIMD math kernel achieves approximately 50 nanoseconds per 768-dimensional cosine similarity comparison on ARM64. Our implementation leverages compiler-assisted NEON auto-vectorization (via AppleClang -O3 -ffast-math) and explicit SIMDe intrinsics, approximately 4,300 faster than interpreted pure Python loops and 30 faster than NumPy backed by Accelerate BLAS.
Analysis. These gains stem from three synergistic factors. First, the SIMDe library transparently maps AVX-512 intrinsics to ARM NEON instructions at compile time, enabling portable SIMD code across x86-64 and ARM64 architectures. Second, modern compilers (AppleClang 17+) with -O3 -ffast-math auto-vectorize scalar loops to equivalent NEON throughput, validating our intrinsic-based implementation against the compiler baseline. Third, explicit 4 loop unrolling maximizes instruction-level parallelism by keeping multiple vector registers in flight simultaneously.
The practical implication is substantial: at 50ns per comparison, the math kernel alone can evaluate 20 million vector pairs per second on a single core. This raw throughput is the engine that powers the SLB’s brute-force strategy—scanning 64 cached nodes takes approximately 3.4 microseconds, well within L1/L2 cache access latency budgets.
@@ 6.2 Macro-Benchmarks: The SLB Impact
Hit latency: 5s (measured 3.7s). The SLB scan executes entirely within the L1/L2 cache hierarchy, requiring only a single SIMD pass over 64 cached centroids.
Miss latency: 10.7s. The search falls back to a root-to-leaf traversal of the Atlas tree, which at depth 3–4 requires approximately 210 vector comparisons at 50ns each.
@@
Comparison. The HNSW baseline exhibits constant latency of approximately 1.5ms regardless of query locality, as it cannot exploit temporal correlation between queries. Consequently, Aeon outperforms HNSW by over 300 on average (s vs. 1.5ms), and by over 400 on cache hits. This advantage grows with workload locality: in highly focused conversational sessions, hit rates approach 95%, reducing effective latency to under 4s.
@@ 6.3 Scalability: From 10K to 1M Nodes
Result. Flat (brute-force) search exhibits linear scaling: latency grows proportionally with database size, reaching 72ms at nodes. In contrast, Aeon Atlas demonstrates logarithmic scaling: traversal latency increases from 7.1s (10K nodes, depth 2) to 10.7s (100K nodes, depth 3) and plateaus at 10.7s (1M nodes, depth 4). This represents an acceleration of over three orders of magnitude (6,000) against flat scan at one million nodes.
Analysis. This behavior is a direct consequence of the B+ tree structure underlying the Atlas. Each level of the tree partitions the search space by a branching factor of , yielding complexity for tree traversal. At nodes, the tree depth is only levels, requiring approximately cosine similarity comparisons. At 50ns per comparison, the theoretical lower bound is 12.8s; the measured 10.7s confirms that the traversal executes entirely within the L1/L2 cache hierarchy, bypassing traditional database I/O bottlenecks.
@@ 6.4 The Zero-Copy Overhead
Result. Transferring 10MB of vector data from C++ to Python incurs sub-microsecond latency (334ns) when using zero-copy shared memory via nanobind. Traditional serialization of standard Python list[float] payloads, representative of unoptimized “Flat RAG” architectures, imposes severe overhead due to object boxing and heap fragmentation:
JSON serialization: 318ms (10 slower)
Pickle serialization: 32.3ms (10 slower)
Analysis. The zero-copy design ensures that Python can access search results without any data movement. The 334ns overhead consists solely of constructing the NumPy array header and validating memory alignment; no actual bytes are copied. This is achieved through the nanobind library’s ability to expose raw memory pointers as read-only NumPy arrays. The five-order-of-magnitude gap (10) between zero-copy and traditional serialization reflects the fundamental cost of Python object boxing: each float in a standard list consumes 28 bytes of heap-allocated storage versus 4 bytes in a contiguous C array.
Figure 4 visualizes this disparity on a logarithmic scale, illustrating the five-order-of-magnitude gap between zero-copy and traditional Python serialization formats.
The practical implication is profound: the Python Shell can orchestrate complex reasoning chains, invoking the Atlas dozens of times per conversation turn, without incurring serialization penalties. A system using JSON-based IPC would spend more time marshalling data than performing actual computation. Aeon’s architecture eliminates this bottleneck entirely, enabling true real-time performance for agentic applications.
@@ 6.5 Summary
Cache Efficiency: The SLB achieves 85%+ hit rates under realistic workloads, reducing effective latency to 4.75s, over 300 faster than HNSW.
@@
Scalability: Logarithmic scaling ensures 10.7s traversal latency even at 1 million nodes, an acceleration of 6,000 over flat scan.
@@
Integration Overhead: Zero-copy memory sharing eliminates serialization costs, achieving 334ns cross-language transfer.
@@ 7 Conclusion
This paper introduced Aeon, a Cognitive Operating System designed to address the fundamental limitations of stateless Large Language Models in long-horizon agentic contexts. The central argument is that the prevailing view of LLM memory as a simple retrieval problem is insufficient; instead, it must be treated as an active resource management task, governed by principles analogous to those found in classical operating system kernels. By formalizing semantic memory management, this work demonstrates that the chaotic, probabilistic nature of vector search can be transformed into a deterministic, navigable process.
@@ 7.1 Key Contributions
This work makes three primary contributions. First, an architecture built upon the Core-Shell Zero-Copy model was presented. The tight integration of a high-performance C++23 kernel (the Atlas) with a flexible Python reasoning layer (the Trace) via nanobind and shared memory proves viable for achieving both the latency requirements of real-time interaction and the expressiveness needed for complex neuro-symbolic reasoning. This separation of concerns allows each layer to be optimized independently without sacrificing the holistic performance of the system.
Second, it was demonstrated that the Semantic Lookaside Buffer (SLB) reduces effective retrieval latency by over two orders of magnitude for conversational workloads. By exploiting the inherent “semantic inertia” of human dialogue (the empirical observation that consecutive topics are highly correlated), semantically likely nodes are pre-positioned in a small, L1/L2 cache-resident structure. Evaluation showed consistent hit rates exceeding 85% under realistic access patterns, validating the core hypothesis that semantic locality is a predictable and exploitable property.
Third, the Trace graph provides a level of interpretability and control that is absent in purely neural approaches. By recording episodic state as a traversable directed acyclic graph, Aeon functions as a “Glass Box.” Every decision can be traced back through its causal lineage. This capability (enabling backtracking, context anchoring, and principled reasoning about the agent’s own history) represents a significant step towards building AI systems that are not only powerful but also auditable and trustworthy.
@@ 7.2 Limitations
Several limitations in the current design are acknowledged. The Atlas currently relies on static embedding models (e.g., BERT-based encoders). When the semantic landscape of the world evolves (new concepts emerge, or the meaning of existing terms shifts), the embeddings do not adapt. While the Delta Buffer provides a mechanism for ingesting new knowledge without a full rebuild, the underlying semantic geometry remains fixed. Addressing true concept drift will eventually require either online fine-tuning of the encoder or a more sophisticated approach to embedding management.
Furthermore, Aeon is presently a single-modality system, operating exclusively on text. Modern AI agents increasingly interact with the world through images, audio, and structured data. A truly general Cognitive OS must eventually support multimodal vector representations within the same spatial index, enabling unified retrieval across diverse input streams, an avenue left for future investigation.
@@ 7.3 Future Work
Aeon represents a first step toward a new paradigm in AI memory management, and several exciting directions remain. Future work envisions Multi-Tenancy and Hardware-Enforced Isolation. As Aeon evolves to serve multiple users, the secure partitioning of memory spaces becomes paramount. The use of hardware enclaves, such as Intel SGX or ARM CCA, is being investigated to provide cryptographic guarantees of data isolation at the memory level.
A “Dreaming” Process is also proposed: an offline background task that activates during idle periods. This process would perform garbage collection on the Atlas, defragmenting the spatial index, and consolidate the verbose Trace into compressed, long-term episodic summaries. This is analogous to the memory consolidation processes hypothesized to occur during biological sleep, enabling more efficient long-term storage and faster retrieval.
Finally, the aim is to integrate Neuro-Symbolic Reasoning more deeply into the architecture. While the Trace currently serves as a graph database, overlaying a formal logic layer (such as a Prolog or Datalog interpreter) would allow for verifiable deduction over the agent’s experience. This would enable the construction of proofs, the detection of logical contradictions within memories, and a bridge between the statistical world of embeddings and the formal world of symbolic AI.
@@ 7.4 Closing Remarks
The path forward for AI agents lies not in ever-larger context windows, but in smarter, more structured memory. Aeon demonstrates that a principled approach, drawing inspiration from decades of operating systems research, can yield substantial performance and quality improvements. This work aims to contribute a useful foundation for researchers and practitioners seeking to build the next generation of persistent, coherent, and interpretable AI agents.